OpenStudy (anonymous):

need help on 3^x-4=9√3

Directrix (directrix):

What are the instructions?

OpenStudy (anonymous):

Solve for x. \[3^x-4=9√3\] Add 4 to both sides: \[3^x=9√3+4\] Make it a logarithm: \[\log_{3}9\sqrt3+4=x \] Plug into your calculator: x=2.71